OpenStudy (anonymous):

Use a table to solve. Round to the nearest hundredth. 2^8x = 93 0.82 30.72 3.14 52.31

OpenStudy (anonymous):

a table? what is this, 1957?

OpenStudy (anonymous):

\[2^{8x}=93\] \[8x=\frac{\ln(93)}{\ln(2)}\] \[x=\frac{\ln(93)}{8\ln(2)}\]

OpenStudy (anonymous):

not sure what "table' you are supposed to use (a log table?) but it looks like around \(.82\) http://www.wolframalpha.com/input/?i=ln%2893%29%2F%288ln%282%29%29

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

yw